The trade balance, B X-M where X Exports, M- imports Show, in details, that the variance of B V(X)V(M) -2Cov(X,M)
Answer #1

Let B=X-M

Var(B)= E[ (B – E(B)2 ]     

= E(B2) - B2 where B2 is the expected value of E(B)

=E[ (X-\small \muX)2 - 2(M-\small \muM)(X-\small \muX) + (M-\small \muM)2 ]  

= Var (X) - 2Cov(X,M) + Var(M)

Since we know Var(X-Y)= Var(X) + Var(Y) - 2Cov(X,Y)

and  \small \muX= E(X) and  \small \muM=E(M)